Please answer all the questions using Excel.

Dean takes his boat out fishing every weekend. His current boat is still in okay condition, but he decides he’d like to buy a new one. He finds the boat of his dreams for $20,950. He does some research and finds that his credit union will give him a 5-year loan with an APR of 4.25% if he makes a down payment of 18%.

  1. What is the amount of Dean’s down-payment?   
  1. What is the loan amount that Dean will be financing?

  1. What will his monthly payments be on the loan?
  1. What is the total amount Dean will pay for his boat if he finances the purchase?
  1. How much interest will Dean pay the credit union over the life of the loan?   

Please, show me your all works. Thanks.
Answer #1

Down Payment = 0.18(20,950) = $3,771

Loan Amount = (1 - 0.18)(20.950) = $17,179

Calculating Monthly Payment,

Using TVM Calculation,

PMT = [PV = 17,179, FV = 0, N = 60, I = 0.0425/12]

PMT = $318.32

Monthly Payment = $318.32

Total Amount Paid = 60(318.32)

Total Amount Paid = $19,099.20